# On estimation and order selection for multivariate extremes via clustering

### v2

@@ 6.1 Simulation studies
In the supplementary material [5], we include more simulation studies as above with different choices of extremal subsample sizes. The results are similar in terms of the bias correction effect of the penalty. We also observe that when the extremal subsample size is chosen relatively large (e.g. $1500$), the accuracy of order selection deteriorates, which is expected since the effect of the convergence of Corollary 2.6 is not yet strong with a low threshold.
@@ 6.2.1 Air Pollution Data
Furthermore, Figures 7 and 9 include visualizations of cluster centers computed based on the $k$-pc algorithm of [10] for the two datasets when we choose the numbers of clusters as above, respectively. Each row in either of the plots corresponds to the coordinate vector of a cluster center: a deeper shade of color indicates a higher value of the squared coordinate. Note that since we work with the $2$-norm sphere, the squared coordinates for each cluster center sum up to $1$, forming a probability distribution row-wise. For the summer data in Figure 7, whose order has been chosen as $5$, the cluster centers concentrate sharply near coordinate directions, which to an extent indicates an asymptotic (or say extremal) independence (see, e.g., [1, Chapter 8]) of the pollutants. In contrast, for the winter data in Figure 9, whose order has been chosen as $3$, a cluster center indicates a group of concomitant extremes consisting of NO, NO2 and PM10. The asymptotic dependence between these 3 variables has been observed in [17]. This serves as a support for our order choice which has placed these 3 variables in the same concomitant group.
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Figures 7 and 9.
@@ 6.2.2 River Discharge Data
BELOW SPRAGUE RIVER NEAR CHILOQUIN, OR WILLIAMSON RIVER 2
@@
As in the previous example, Figure 12 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Figure 12 illustrates the squared cluster centers obtained from the $k$-pc algorithm when the order is chosen as $6$. In Figure 12 , we convert the spectral estimation to the factor matrix $B$ following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Figure 12) the largest value (in bold) corresponds. We include these factor indices in the last column of Table 1, which can be viewed roughly as markings of groups of concomitant extremes. The results align well with the expected geographical context: These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in Minnesota (MN) and the other located in Illinois (IL).
@@ 7 Summary and Discussion
At last, we point out several potential future work directions. First, the tuning parameter $t$ for the penalty (22) is chosen with the help of visual inspection. It is of interest to explore the data-driven method to guide the choice of $t$, which may require a refined understanding of the consistency result in Theorem 3.1. In addition, it may be worth exploring alternative clustering assessment criteria to ASW in (21). Some preliminary experiments show that the cross-validation method based on algorithmic instability of [32] might also be a competent criterion in the context of clustering multivariate extremes. It would also be desirable to develop a method for choosing the threshold $l_{n}$ in (14) in the context of clustering multivariate extremes. The recent work [31] might be relevant in this regard. Furthermore, it may be of interest to enhance the large-deviation-type analysis in Section 4 to a finite-sample probabilistic guarantee bound.

### v3

@@ 6.1 Simulation studies
In the supplementary material [5], we include more simulation studies as above with different choices of extremal subsample sizes. The results are similar in terms of the bias correction effect of the penalty. We also observe that when the extremal subsample size is chosen relatively large (e.g., $1500$), the accuracy of order selection deteriorates, which is expected since the effect of the convergence of Corollary 1 is not yet strong with a low threshold.
@@ 6.2.1 Air Pollution Data
Following the method introduced in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$, we compute the factor coefficient matrix $B$ for the two datasets; see Figures 7 and 9. Furthermore, Figures 7 and 9 include visualizations of cluster centers computed based on the $k$-pc algorithm of [10] for the two datasets when we choose the numbers of clusters as above, respectively. Each row in either of the plots corresponds to the coordinate vector of a cluster center: a deeper shade of color indicates a higher value of the coordinate. For the summer data in Figure 7, whose order has been chosen as $5$, the cluster centers concentrate sharply near coordinate directions, which to an extent indicates an asymptotic (or say extremal) independence (see, e.g., [1, Chapter 8]) of the pollutants. In contrast, for the winter data in Figure 9, whose order has been chosen as $3$, a cluster center indicates a group of concomitant extremes consisting of NO, NO2 and PM10. The asymptotic dependence between these 3 variables has been observed in [17]. This serves as a support for our order choice which has placed these 3 variables in the same concomitant group.
@@ 6.2.2 River Discharge Data
BELOW SPRAGUE RIVER NEAR CHILOQUIN, OR WILLIAMSON RIVER 2
@@
As in the previous example, Figure 12 presents the penalized ASW curves, from which we found that 6 seems to be an appropriate choice of order. Figure 12 illustrates the cluster centers obtained from the $k$-pc algorithm when the order is chosen as $6$. In Figure 12 , we convert the spectral estimation to the factor matrix $B$ following the method in Section 5.2 with $\|\cdot\|_{(s)}=\|\cdot\|_{(r)}=\|\cdot\|_{2}$ and $\alpha=2$. In addition, for each row of the matrix $B$, we find to which factor index (the same as the cluster index in Figure 12) the largest value corresponds. We include these factor indices in the last column of Table 1, which can be viewed roughly as markings of groups of concomitant extremes. The results align well with the expected geographical context: These 6 groups are in good accordance with the geographical context: most of the stations located along the same river are found in the same group, with the only exception of the 4 stations along the Mississippi River. The further division of these 4 stations into 2 groups may be easily justified by the large geographical distance between the 2 groups: one group located in Minnesota (MN) and the other located in Illinois (IL).
@@ 7 Summary and Discussion
At last, we point out several potential future work directions. First, the tuning parameter $t$ for the penalty (22) is chosen with the help of visual inspection. It is of interest to explore the data-driven method to guide the choice of $t$, which may require a refined understanding of the consistency result in Theorem 1. In addition, it may be worth exploring alternative clustering assessment criteria to ASW in (21). Some preliminary experiments show that the cross-validation method based on algorithmic instability of [32] might also be a competent criterion in the context of clustering multivariate extremes. It would also be desirable to develop a method for choosing the threshold $l_{n}$ in (14) in the context of clustering multivariate extremes. The recent work [31] might be relevant in this regard. Furthermore, it may be of interest to enhance the large-deviation-type analysis in Section 4 to a finite-sample probabilistic guarantee bound.